es Perform the calculation. 63°26 -23°49 BE FEB 63°26 -23°49 =
Answer #1

63° 26 - 23° 49 Salution * = 60) 63° 26 = 62 (6&t 26) 163 26 62° 86 Therefore, 63° 26-23°49 = 62° 86- 23°49 16828=62